Aaron Abbot invested P100,000 at 12% interest, compounded semi-annually, for 3 years. What is the future value?


1
Expert's answer
2021-09-24T13:15:29-0400

"I = {P(1+\frac{r}{n})}^{nt}"

"Year1 = {100,000\ *\ (1+\frac{0.12}{2})}^{(2*1)} = 112360"


"Year2 = {112,360\ *\ (1+\frac{0.12}{2})}^{(2*1)} = 126,247.70"


"Year3 = {126,247.70\ *\ (1+\frac{0.12}{2})}^{(2*1)} = 141,851.91"


"1st\ year\ interest \ will\ be\ \$12,360"


"2nd\ year\ interest \ will\ be\ \$13,887.70"


"3rd\ year\ interest \ will\ be\ \$15,604.21"


"the \ total\ interest\ earned\ will\ be \ \$ 41,851.91\ after\ 3years"


"the \ compound\ amount \ earned\ will\ be \ \$ 141,851.91\ after\ 3years"